H and F got married 21 ye ago F was 3 times older then H today F is only twice as old how old was H 21 ye ago
Answer by Cromlix(4381)

21 years ago F was 3 times older then H
F = 3H
Today F is only twice as old as H
F + 21 = 2(H + 21)
Replace F with 3H
3H + 21 = 2(H + 21)
3H + 21 = 2H + 42
3H - 2H = 42 - 21
H = 21 years old.
Hope this helps.
:-)